OpenStudy (anonymous):

Simplify: x^2 - 16/4x divided by ( x-4 )

OpenStudy (callisto):

Apply the identity \(a^2 - b^2 = (a+b)(a-b)\) to factor x^2 - 16 \[\large \frac{\frac{x^2 - 16}{4x}}{( x-4 )}\]\[\large =\frac{\frac{(x+4)(x-4)}{4x}}{( x-4 )}\]\[\large =\frac{(x+4)(x-4)}{4x( x-4 )}\]Cancel the common factor. Can you do it?

OpenStudy (anonymous):

yeah its x+4/4x

OpenStudy (callisto):

Yup :)

OpenStudy (anonymous):

thanks for helping :)